# Joint Score-Threshold Optimization for Interpretable Risk Assessment Under Partial Supervision

### v1

@@ 3.2 MIP Models
The JHFRAT scale is divided, per current practice, into three risk categories: Low-Risk ($k=1$) for scores less than 6, Moderate-Risk ($k=2$) for scores 6-13, and High-Risk ($k=3$) for scores greater than 13. While our general framework jointly optimizes both scoring weights $\beta$ and thresholds $\tau$, the JHFRAT application presents unique deployment constraints that motivate a modified approach. To maintain compatibility with existing clinical workflows, electronic health record systems, and staff training, we fix the thresholds at their current values ($\tau_{1}=6,\tau_{2}=13$) and focus optimization on the scoring weights $\beta$ alone. This decision reflects a common practical scenario where healthcare systems require incremental refinements to existing tools rather than complete redesigns, as changing risk thresholds would necessitate updates to intervention protocols, clinical guidelines, and quality metrics across the entire health system.
The absence of reliable moderate-risk labels in our dataset further supports this constrained optimization approach. Due to the selective labeling problem, we can confidently identify only extreme cases: patients who fell despite interventions (Safe High-Risk) and those who remained fall-free without targeted interventions (Safe Low-Risk). Patients with moderate JHFRAT scores predominantly received interventions (78.3% in our data), making their true risk indeterminate—they might be correctly classified moderate-risk patients, or they might be misclassified high or low-risk patients. Rather than introducing noise by guessing labels for this ambiguous middle group, we apply singleton feasible sets only to the extreme cases: $\mathcal{S}_{i}=\{1\}$ for all $i$ in the Safe Low-Risk cohort and $\mathcal{S}_{i}=\{3\}$ for all $i$ in the Safe High-Risk cohort.
This configuration represents a special but important case of our general framework where partial supervision is extreme—no labeled examples exist for intermediate categories. The fixed thresholds naturally prevent the middle category collapse problem described in Section 2.3, as the moderate-risk band maintains its width of $\tau_{2}-\tau_{1}=7$ points regardless of the optimization. While this approach doesn’t leverage our framework’s full capability for joint score-threshold learning, it demonstrates its flexibility in handling severely limited supervision scenarios common in healthcare, where intervention effects obscure risk for the majority of patients. Future work could explore threshold optimization using external validation data or prospective trials where moderate-risk labels can be more reliably determined.
We apply three types of governance constraints in our MIP formulations for JHFRAT. Firstly, we apply a uniform lower and upper bound on all model coefficients. Since we only include the risk factors from JHFRAT with positive coefficients, it is reasonable to impose the clinical assumption inherent in the lower bound of 0 that all of these factors can only contribute positively to high fall risk. Furthermore, we limit coefficients to the highest class’s lower bound of 13 for model stability. Secondly, we impose integer constraints on the coefficients to maintain consistency with the format of JHFRAT and ease cognitive burden of manual assessment. Finally, we impose monotonicity constraints for hierarchical risk factors to mantain clinical validity. In particular, the risk factors in the categories of age, medications, and patient care equipment are inherently hierarchical, and we let $P$ denote the set of pairs of risk factor indices corresponding to these pairwise orderings (i.e. the model coefficient for one high-fall risk drug is less than or equal to the coefficient for two or more high-fall risk drugs).
The complete JHFRAT-optimizing MIP formulation is:
$\min_{\beta,\tau,y,z}\quad$ $\sum_{i=1}^{n}w_{i}\sum_{k=1}^{3}c(k,\mathcal{S}_{i})\cdot z_{ik}$ (19a)
@@
(1d) - (1h) (19c)
$0\leq\beta_{j}\leq 13\quad\forall j$ (19d)
$\beta_{i}\leq\beta_{j}\quad\forall(i,j)\in P$ (19e)
$\beta_{j}\in\mathbb{Z}\quad y_{ik},z_{ik}\in\{0,1\}$ (19f)
We consider four variations to the MIP model above:
  1. 1.
    Symmetric: The ordinal loss for under-triage and over-triage is the same ($\alpha_{under}=\alpha_{over}=1$). Constraints and other hyperparameters follow exactly from Model (19).
  2. 2.
    Asymmetric: The ordinal loss for under-triage is three times that of over-triage ($\alpha_{under}=3,\alpha_{over}=1$). Constraints and other hyperparameters follow exactly from Model (19).
  3. 3.
    False-Positive (FP) constrained: The ordinal loss for under-triage is three times that of over-triage, and all other hyperparameters follow exactly from Model (19). Let $\bar{\beta}$ indicate the JHFRAT model coefficients and let $\bar{z}$ indicate the feasible set binary classification variables imposed deterministically by the MIP formulation with $\bar{\beta}$. We impose the following two constraints, in addition to those in Model (19):
    $\sum_{i\in[n]|0\in S_{i}}z_{i1}$ $\leq\sum_{i\in[n]|0\in S_{i}}\bar{z}_{i1}$ (20a)
    $\sum_{i\in[n]|0\in S_{i}}z_{i2}$ $\leq\sum_{i\in[n]|0\in S_{i}}\bar{z}_{i2}$ (20b)
@@
    These additional performance constraints ensure that the number of Safe Low-Risk labeled encounters classified as Moderate-Risk (20a) and High-Risk (20b) cannot exceed the numbers of such classifications with the JHFRAT model.
  4. 4.
    False-Negative (FN) & False-Positive (FP) constrained: The ordinal loss for under-triage is three times that of over-triage. The ordinal loss for under-triage is three times that of over-triage, and all other hyperparameters follow exactly from Model (19). With the same definition of $\bar{z}$ as above, we impose the following four constraints, in addition to those in Model (19):
    $\sum_{i\in[n]|2\in S_{i}}z_{i0}$ $\leq\sum_{i\in[n]|2\in S_{i}}\bar{z}_{i0}$ (21a)
    $\sum_{i\in[n]|2\in S_{i}}z_{i1}$ $\leq\sum_{i\in[n]|2\in S_{i}}\bar{z}_{i1}$ (21b)
    $\sum_{i\in[n]|0\in S_{i}}z_{i1}$ $\leq 0.05\sum_{i\in[n]|0\in S_{i}}1$ (21c)
    $\sum_{i\in[n]|0\in S_{i}}z_{i2}$ $\leq 0.3\sum_{i\in[n]|0\in S_{i}}1$ (21d)
    The number of Safe High-Risk labeled encounters classified as Low-Risk (21a) and Moderate Risk (21b) cannot exceed the numbers of such classifications with the JHFRAT model. The portion of Safe Low-Risk labeled encounters classified as Moderate-Risk (21c) and High-Risk (21d) cannot exceed given thresholds (30% and 5% respectively). These constraints allow for a small increase from the JHFRAT benchmark of 26.5% and 0.9% for these same categorizations, as seen in Table 2.
We note that this fixed-threshold approach, while more restrictive than full joint optimization, offers several practical advantages: (1) it ensures the optimized tool remains immediately deployable without systemic changes to care protocols, (2) it maintains comparability with historical JHFRAT data and quality metrics, and (3) it reduces the optimization complexity, improving solution stability and interpretability. The resulting problem focuses on finding optimal integer weights $\beta\in\mathbb{Z}^{17}_{+}$ that best separate the extreme risk groups within the established scoring scale, effectively treating this as a constrained parameter refinement problem rather than a full model redesign.
@@ 3.3 Experimental Setup
We split the encounters from the combined fall and reference cohorts into an 80% (encounters) training and validation set, and a 20% (encounters) test set, stratified by risk label. We retain the intervention cohort separately as an exogenous dataset for concordance analysis post-optimization. We perform 5-fold cross-validation for training the CSO, and run a hyperparameter gridsearch over the values of $\alpha_{under}$, $\alpha_{over}\in\{0.5,1,2,4\}$ and $\lambda_{1}\in\{0.2,0.5,0.8\}$ (such that $\lambda_{1}=\lambda_{1}^{-}=\lambda_{1}^{+}$ and $\lambda_{2}=1-\lambda_{1}=\lambda_{2}^{-}=\lambda_{2}^{+}$). For each set of hyperparameters, we consider the average coefficients across the cross-validation folds. Each MIP model is trained on the entire training set, utilizing the best feasible rounded coefficients from the CSO gridsearch. If none of the rounded CSO coefficients are feasible for a given MIP model, the JHFRAT coefficients are used for the warm-start.The CSO model is solved with cvxpy, and all MIP variations are solved with Gurobi 12.0.3.
@@ 3.4 Baselines and Evaluation Metrics
We compare the outcomes of our approach against the original JHFRAT model with expert-derived weights. Furthermore, we compare the CSO performance with those of the MIP models in order to evaluate the added value of the MIP portion of the optimization pipeline. We utilize the following metrics to compare model performance:
  • Classification Accuracy: We define tight accuracy as the portion of encounters correctly classified into low or high-risk categories. We further define loose accuracy to additionally include all encounters classified as moderate-risk as correctly classified.
  • High-Risk Precision and Recall: To account for the imbalance of the dataset, we rely on the high-risk precision and recall in addition to the accuracy metrics.
  • Binary Classification Performance Measures: We utilize the distribution of scores for the optimized models to generate receiver operating characteristic (ROC) and precision-recall curves. We then utilize the area under each curve, AUROC and AUPRC respectively, as model performance metrics.
@@ 4 Results
4.1 Optimized Risk Factor Coefficients
The optimal solutions to the four MIP variations are featured in Table 2, in comparison with the current JHFRAT scoring coefficients. The optimal CSO coefficients for symmetric class loss (hyperparameters $\alpha_{under},\alpha_{over}=1,\lambda_{1}^{-},\lambda_{1}^{+},\lambda_{2}^{-},\lambda_{2}^{+}=1$) are also included as a representative example of CSO results.
@@ 4.1 Optimized Risk Factor Coefficients
Risk Factor JHFRAT CSO Symmetric Asymmetric FP-cap FN&FP-cap
Age 60-69 years 1 1.70 1 0 1 0
Age 70-79 years 2 1.92 2 3 1 1
Age $\geq$ 80 years 3 2.22 2 13 4 2
Incontinence 2 7.33 11 9 4 12
Urgency/frequency 2 3.75 3 6 5 3
Altered awareness 1 7.66 12 9 12 10
Impulsive 2 15.01 13 11 12 7
Lack of understanding 4 5.20 13 12 3 12
One present 1 0.77 0 3 2 2
Two present 2 1.38 1 3 3 6
Three or more present 3 1.56 2 13 3 8
Fall within 6 months 5 2.22 0 6 0 0
One HFRD 3 2.60 2 2 1 0
Two or more HFRD 5 4.04 4 3 3 0
Sedation procedure 7 4.04 4 3 3 0
Requires assistance 2 6.84 11 13 5 8
Unsteady gait 2 6.76 13 10 0 7
Visual/auditory impairment 2 6.92 6 2 9 11
Total Coefficient Sum 49 81.92 100 121 71 89
All optimized models have total coefficient sums greater than the current JHFRAT, and thus we compare feature contributions via their percent contributions to the total coefficient sum. A few trends persist across the model variations. Cognition risk factors feature far more prominently in the MIP models than in the current JHFRAT. Conversely, the history of falls risk factor only persists as non-zero in the asymmetric MIP model, albeit with less proportional importance (7.4%) compared to JHFRAT (10.2%).
4.2 Score Distributions and Categorization
Figure 2 reveals how different optimization objectives reshape the risk score distributions compared to the original JHFRAT. The Kolmogorov-Smirnov (KS) statistic quantifies the maximum vertical separation between the cumulative distribution functions of the Safe Low-Risk and Safe High-Risk cohorts, with larger values indicating better discriminative ability. The associated p-values test the null hypothesis that the two risk groups come from the same distribution—highly significant p-values (all p < 0.001) confirm that all models successfully separate the risk groups, though with varying degrees of separation.
Figure 2: Score distributions for all labeled data (test and train data together)
The distribution patterns directly reflect each model’s clinical optimization strategy:
Original JHFRAT–Conservative but Imprecise: The overlapping distributions (KS = 0.620) show why current practice struggles: many true high-risk patients score in the low-to-moderate range (under-triage), while some low-risk patients score high (over-triage). This overlap drives both missed falls and unnecessary interventions. With 44.8% of eventual fall patients classified as low or moderate risk, critical prevention opportunities are missed.
Symmetric MIP–Maximum Separation for Clear Decision-Making: The stark bimodal distribution (KS = 0.699) eliminates the ambiguous middle ground, pushing patients toward clear low-risk or high-risk classifications. This suits healthcare systems that prefer definitive decisions: either minimal monitoring (score < 6) or full prevention protocols (score > 13). The trade-off is extreme scores (up to 60) that may seem implausible to clinicians, potentially undermining trust. Best for: facilities with binary intervention protocols where middle-ground treatments offer little value.
Asymmetric MIP–Aggressive Fall Prevention: The rightward shift of both distributions reflects the 3:1 penalty for under-triage, effectively lowering the bar for high-risk classification. This achieves 86% sensitivity for falls (vs. 45% in original JHFRAT) but classifies 17% of safe low-risk patients as high-risk. The extreme score range (up to 70) results from the optimizer pushing borderline cases strongly rightward to avoid missing falls. Best for: facilities where fall consequences are severe (e.g., surgical or elderly units) and resources allow broader intervention deployment.
FP-cap MIP–Targeted High-Risk Identification: By constraining false positives to current levels, this model (KS = 0.671) maintains clinician workload while better identifying true high-risk patients. The distribution shows selective rightward movement—pushing true fall patients across the high-risk threshold while keeping most low-risk patients in place. Scores remain clinically plausible (maximum 35). This achieves 45% fall sensitivity with 81% positive predictive value, meaning when it flags high-risk, it’s usually correct. Best for: resource-constrained settings where intervention capacity is fixed and false alarms cause alert fatigue.
FP&FN-cap MIP—Balanced Improvement Within Constraints: The bidirectional constraints produce the most JHFRAT-like distribution while still achieving better separation (KS = 0.683 vs. 0.620). Both cohorts shift toward the extremes but within bounds that maintain operational compatibility. This model essentially "tune” JHFRAT rather than replacing it, achieving 64% fall sensitivity while keeping false positive rates manageable. Best for: conservative healthcare systems requiring evidence that new models won’t disrupt operations before accepting larger changes.
CS–Smooth Risk Stratification: The continuous optimization produces graduated separation without the stark bimodality of MIP solutions. The smooth distribution (KS = 0.700) suggests CSO finds a natural risk continuum rather than forcing binary separation. This may better reflect clinical reality where risk truly is continuous, not discrete. With 66% fall sensitivity and moderate false positives, it balances competing objectives. Best for: initial implementations where extreme changes might face resistance, or as a warm-start for subsequent MIP refinement.
The key insight is that no single distribution is universally "best"–each reflects different clinical priorities and operational constraints. The original JHFRAT’s overlapping distributions explain its poor performance, while our optimized models offer a menu of alternatives: maximum separation (Symmetric MIP) for clear decisions, protective bias (Asymmetric MIP) for high-stakes environments, resource-aware precision (FP-cap) for efficiency, or incremental improvement (FP&FN-cap) for conservative adoption. The extreme scores in unconstrained models, while mathematically optimal, highlight the importance of including clinical face-validity constraints in future deployments.
We calculate the score differentials (MIP score - JHFRAT score) for all encounters in the dataset, as seen in Figure 3. All MIP models results in a positive average score differential, with a heavy upper tail. The asymmetric MIP without any performance constraints results in the most positive skew and the FP-cap MIP results in the least positive skew.
Figure 3: Score differentials (MIP score - JHFRAT score) for all patients in dataset
Skewness in the score differentials is also reflected in the categorization of encounters across the low, moderate and high-risk categories shown in Table 3. Without additional performance constraints, both the symmetric and asymmetric models significantly increase the number of encounters categorized as high-risk in the test data set. The symmetric model results in a 1100% increase in high-risk false positives and a 164% increase in true positives while the asymmetric model has a 1878% increase in high-risk false positives and 200% increase in true positives in the test set. The two models with false positive maximums still achieve notable gains in true positives: 56% and 124% increases from JHFRAT for FP-cap and FN&FP-cap, respectively. Both models also decrease the number of safe high-risk patients categorized as moderate-risk, but increase the number categorized as low-risk. All models except for the asymmetric MIP without performance constraints achieve a small increase in the number of true negative safe low-risk patients identified as low-risk.
@@ 4.2 Score Distributions and Categorization
Table 3: Test data encounter categorization across baseline and optimized models
Model Encounters Model Categorization
Low-Risk Moderate-Risk High-Risk
Original JHFRAT Safe Low-Risk 1,149 (72.6%) 420 (26.5%) 14 (0.9%)
Safe High-Risk 7 (8.1%) 55 (63.2%) 25 (28.7%)
All 17,007 (31.4%) 30,254 (55.8%) 6,946 (12.8%)
CSO Safe Low-Risk 1,156 (73.0%) 374 (20.5%) 53 (3.3%)
Safe High-Risk 8 (9.2%) 22 (25.3%) 57 (65.5%)
All 15,640 (28.9%) 18,124 (33.43%) 20,443 (37.7%)
Symmetric MIP Safe Low-Risk 1,180 (74.5%) 262 (16.6%) 168 (10.6%)
Safe High-Risk 7 (8.1%) 14 (16.1%) 66 (75.9%)
All 16,207 (29.9%) 10,581 (19.5%) 27,419 (50.58%)
Asymmetric MIP Safe Low-Risk 946 (59.8%) 360 (22.7%) 277 (17.5%)
Safe High-Risk 6 (6.9%) 6 (6.9%) 75 (86.2%)
All 11,278 (20.8%) 9,212 (17.0%) 33,717 (62.2%)
FP-cap MIP Safe Low-Risk 1,258 (79.5%) 316 (20.0%) 9 (0.6%)
Safe High-Risk 12 (13.8%) 36 (41.4%) 39 (44.8%)
All 18,142 (33.5%) 22,892 (42.2%) 13,172 (24.30%)
FP&FN-cap MIP Safe Low-Risk 1,237 (78.1%) 294 (18.6%) 52 (3.3%)
Safe High-Risk 11 (12.6%) 20 (23.0%) 56 (64.4%)
All 18,034 (33.3%) 15,562 (28.7%) 20,611 (38.0%)
4.3 Performance Metrics
Table 4 includes the performance metrics calculated for each model. In general, the optimized models achieve moderate gains in tight accuracy at the expense of loose accuracy. Only the FP-cap MIP improves on the tight accuracy without loss of loose accuracy.
@@ 4.3 Performance Metrics
Model Accuracy AUC High-Risk Precision High-Risk Recall
Tight Loose AUROC AUPRC
Original JHFRAT 0.70 0.99 0.92 0.51 0.64 0.29
CSO 0.73 0.96 0.94 0.65 0.52 0.66
Symmetric MIP 0.75 0.90 0.94 0.65 0.28 0.76
Asymmetric MIP 0.61 0.83 0.92 0.57 0.21 0.86
FP-cap MIP 0.78 0.99 0.92 0.62 0.81 0.45
FN&FP-cap MIP 0.77 0.96 0.90 0.63 0.52 0.64
All optimization models outperform JHFRAT in area under the precision-recall curve for the testing and training datasets. All models perform similarly with high recall in the testing dataset, with the curves only notably separating at recall below 0.6. In contrast, the optimized models achieve visibly higher precision across all recall points for the training data. The small improvements in AUROC achieved by the models in the training set do not materialize in the testing set.
Figure 4: Receiver Operating Characteristic (ROC) and Precision-Recall performance curves across JHFRAT and five optimization models on training dataset (above) and testing dataset (below).

### v2

@@ 4.2 MIP Models
The JHFRAT scale is divided, per current practice, into three risk categories: Low-Risk ($k=1$) for scores less than 6, Moderate-Risk ($k=2$) for scores 6-13, and High-Risk ($k=3$) for scores greater than 13. We apply the corresponding true labels to the labeled training data: $k^{*}=1$ for the low-risk patients in the reference cohort and $k^{*}=3$ for the high-risk patients in the fall cohort. While our general framework jointly optimizes both scoring weights $\beta$ and thresholds $\tau$, the JHFRAT application presents unique deployment constraints that motivate a modified approach. To maintain compatibility with existing clinical workflows, electronic health record systems, and staff training, we fix the thresholds at their current values ($\tau_{1}=6,\tau_{2}=13$) and focus optimization on the scoring weights $\beta$ alone. This decision reflects a common practical scenario where healthcare systems require incremental refinements to existing tools rather than complete redesigns, as changing risk thresholds would necessitate updates to intervention protocols, clinical guidelines, and quality metrics across the entire health system.
We note that this fixed-threshold approach, while more restrictive than full joint optimization, offers several practical advantages: (1) it ensures the optimized tool remains immediately deployable without systemic changes to care protocols, (2) it maintains comparability with historical JHFRAT data and quality metrics, and (3) it reduces the optimization complexity, improving solution stability and interpretability. The resulting problem focuses on finding optimal integer weights $\beta\in\mathbb{Z}^{17}_{+}$ that best separate the extreme risk groups within the established scoring scale, effectively treating this as a constrained parameter refinement problem rather than a full model redesign. The complete JHFRAT-optimizing MIP formulation with these fixed thresholds is given as follows:
$\min_{\beta,\tau,z}\quad$ $\sum_{i=1}^{n}w_{i}\sum_{k=1}^{3}\ell(k,k_{i}^{*})\cdot z_{ik}$ (11a)
@@
(1d) - (1h)
$0\leq\beta_{j}\leq 13\quad\forall j$ (11c)
$\beta_{i}\leq\beta_{j}\quad\forall(i,j)\in P$ (11d)
$\frac{\sum_{i\in[n]:k_{i}^{*}=3}z_{i1}}{|\{i:k^{*}_{i}=3\}|}\leq\xi_{sFNR}$ (11e)
$\frac{\sum_{i\in[n]:k_{i}^{*}=3}z_{i2}}{|\{i:k^{*}_{i}=3\}|}\leq\xi_{mFNR}$ (11f)
$\frac{\sum_{i\in[n]:k_{i}^{*}=1}z_{i3}}{|\{i:k^{*}_{i}=1\}|}\leq\xi_{sFPR}$ (11g)
$\frac{\sum_{i\in[n]:k_{i}^{*}=1}z_{i2}}{|\{i:k^{*}_{i}=1\}|}\leq\xi_{mFPR}$ (11h)
$\beta_{j}\in\mathbb{Z},\quad z_{ik}\in\{0,1\}$ (11i)
The objective function (11a) exactly follows from (1a) by optimizing over all $n$ training samples and 3 JHFRAT categories. The definition for $\ell(k,k_{i}^{*})$ follows from (2), and we set $\omega_{under}=3$ and $\omega_{over}=1$ so that the ordinal loss for under-triage is three times that of over-triage. The sample weights $w_{i}$ are derived from the number of samples in each class such that the total contribution to the objective function is equal between the imbalanced low and high-risk classes: $w_{i}=\frac{n}{|\{j:k_{j}^{*}=k_{i}^{*}\}|}.$
We apply three types of governance constraints in our MIP formulations for JHFRAT. Firstly, we apply a uniform lower and upper bound on all model coefficients via (11c). Since we only include the risk factors from JHFRAT with positive coefficients, it is reasonable to impose the clinical assumption inherent in the lower bound of 0 that all of these factors can only contribute positively to high fall risk. Furthermore, we limit coefficients to the highest class’s lower bound of 13 for model stability. Secondly, in (11d) we impose monotonicity constraints for hierarchical risk factors to maintain clinical validity. In particular, the risk factors in the categories of age, medications, and patient care equipment are inherently hierarchical, and we let $P$ denote the set of pairs of risk factor indices corresponding to these pairwise orderings (i.e., the model coefficient for one high-fall risk drug is less than or equal to the coefficient for two or more high-fall risk drugs). Finally, in (11i) we impose integer constraints on the coefficients to maintain consistency with the format of JHFRAT and ease cognitive burden of manual assessment.
@@
We impose four performance constraints using the number of false positives and false negatives from the original JHFRAT model. The number of High-Risk labeled encounters classified as Low-Risk (11e) and Moderate Risk (11f) cannot exceed the rates of such “severe” and “moderate” false-negative classifications with the JHFRAT model, $\xi_{sFNR}$ and $\xi_{mFPR}$. Conversely, the number of Low-Risk labeled encounters classified as High-Risk (11g) and Moderate Risk (11h) cannot exceed the rates of the severe and moderate false-positive classifications with the JHFRAT model, $\xi_{sFPR}$ and $\xi_{mFPR}$. These performance constraints ensure that improvements achieved by the MIP model do not come at the expense of worse performance for one class or the other.
@@ 4.3 Experimental Setup
We split the encounters from the combined fall and reference cohorts into an 80% training set, and a 20% test set, stratified by risk label and current JHFRAT category. We retain the intervention cohort separately as an exogenous dataset for concordance analysis post-optimization. We first train a series of CSO models via a hyperparameter gridsearch over the values of $\alpha^{-}$, $\alpha^{+}\in\{0.5,1,2,4\}$ and $\lambda_{1}\in\{0.2,0.5,0.8\}$ (such that $\lambda_{1}=\lambda_{1}^{-}=\lambda_{1}^{+}$ and $\lambda_{2}=1-\lambda_{1}=\lambda_{2}^{-}=\lambda_{2}^{+}$). For each set of hyperparameters, we perform 5-fold cross-validation over the training set and average the coefficients across the cross-validation folds. The set of candidate warm-start solutions is the sets of these averaged coefficients, rounded to the nearest integer.
The MIP model is trained on the entire training set, utilizing the candidate warm-start solution that has the best MIP objective value while satisfying all MIP constraints. If no candidate warm-start solutions are feasible, we warm-start with the original JHFRAT coefficients. The CSO models are solved with MOSEK via cvxpy, and all MIP variations are solved with Gurobi 12.0.3.
@@ 4.4 Baselines and Evaluation Metrics
We compare the outcomes of our approach against the original JHFRAT model with expert-derived weights. Furthermore, we compare the CSO performance with that of the MIP models in order to evaluate the added value of the MIP portion of the optimization pipeline. We utilize the following metrics to compare model performance:
4.4.1 Classification Accuracy
An encounter data sample is considered correctly classified only if it is exactly classified into its correct low-risk or high-risk category. A data sample is considered mis-classified with “moderate error” if it is classified into the moderate-risk category, and mis-classified with “severe error” if it is mis-classified into the opposite category: a low-risk labeled sample classified into the high-risk category and vice-versa. We define tight accuracy as the portion of encounters correctly classified, and define loose accuracy to additionally include all instances of moderate error as correctly classified.
4.4.2 High-Risk Precision and Recall
We rely on the high-risk precision and recall to account for the imbalance of the dataset and the importance of correctly identifying high-risk patients in this setting. For these binary-classification-based metrics, moderate and low-risk classifications are grouped together.
4.4.3 Binary Classification Performance Measures
We utilize the distribution of scores for the optimized models to generate receiver operating characteristic (ROC) and precision-recall curves. We then utilize the area under each curve, AUROC and AUPRC, respectively, as model performance metrics.
@@ 5 Results
The optimal solution to the MIP defined in (11) is featured in Table 2, in comparison with the current JHFRAT scoring coefficients. The optimal feasible CSO warm-start coefficients for this MIP (hyperparameters $\alpha^{-}=0.5,\alpha^{+}=1,\lambda_{1}^{-}=\lambda_{1}^{+}=0.8,\lambda_{2}^{-}=\lambda_{2}^{+}=0.2$) are also included. Both the MIP and CSO warm-start have total coefficient sums greater than the current JHFRAT, promoting higher possible total scores. Cognition and elimination risk factors feature far more prominently in the MIP model than in the current JHFRAT. Conversely, all patient care equipment coefficients are zero in the optimal MIP model.
@@
Category Risk Factor JHFRAT CSO MIP
Age 60-69 years 1 1 0
70-79 years 2 1 2
$\geq$ 80 years 3 2 2
Elimination Incontinence 2 6 6
Urgency/frequency 2 3 7
Cognition Altered awareness 1 6 13
Impulsive 2 12 12
Lack of understanding 4 3 4
Equipment 1 present 1 1 0
2 present 2 1 0
3+ present 3 1 0
Fall Hist. Fall within 6 months 5 2 1
Medication 1 high fall risk drug 3 3 2
2+ high fall risk drugs 5 4 2
Sedation procedure 7 4 2
Mobility Requires assistance 2 4 7
Unsteady gait 2 4 2
Visual/auditory impairment 2 6 3
Total Coeff. Sum 49 64 65
Table 3 details the model categorization accuracy per labeled class for each model. Both optimized models improve correct classification on the training data for both low-risk and high-risk patients while reducing or maintaining the rates of moderate and severe error, as designed. For the training data, correct low-risk classification increases from 71% (original JHFRAT) to 83% (MIP), and correct high-risk classification increases from 30% to 46%. This improvement persists in the testing data, with an increases to 84% and 48% correct classification for low-risk and high-risk patients respectively, with the JHFRAT testing data rates remaining the same. Furthermore, the severe and moderate error rates for the models remain the same or lower than the JHFRAT rates for the testing data. The severe error rates remain comparable across models, likely because the severe error performance constraints (5e) and (5g) are in direct conflict with each other: there is a natural limit, to how much the model can reduce the severe false positives without increasing severe false negatives, and vice versa. For the training data, these constraints are explicitly enforced, but the effects of these constraints persist in the testing data, as evident by the consistency in severe error rates for testing data. The optimized models outperform JHFRAT in area under the precision-recall curve for the testing datasets, per Table 4 and Figure 2. The MIP’s focus on categorizing around thresholds is reflected in its improved accuracy metrics over the CSO model, despite having similar AUC scores.
@@
Table 3: Train and test data categorization comparison across original and optimized models
Model Label Categorization
Correct Mod. Err. Severe Err.
Training Data
Orig. JHFRAT Low-Risk 4,519 (71%) 1,788 (28%) 58 (1%)
High-Risk 105 (30%) 202 (58%) $\bm{42(12\%)}$
CSO Low-Risk 4,940 (78%) 1,377 (22%) $\bm{48(1\%)}$
High-Risk 151 (43%) 156 (45%) $\bm{42(12\%)}$
Opt. MIP Low-Risk $\bm{5,271(83\%)}$ $\bm{1,036(16\%)}$ 58 (1%)
High-Risk $\bm{160(46\%)}$ $\bm{147(42\%)}$ $\bm{42(12\%)}$
Testing Data
Orig. JHFRAT Low-Risk 1,135 (71%) 441 (28%) 15 (1%)
High-Risk 26 (30%) 51 (58%) $\bm{11(13\%)}$
CSO Low-Risk 1,279 (80%) 301 (19%) $\bm{11(1\%)}$
High-Risk 39 (45%) 40 (45%) $\bm{9(10\%)}$
Opt. MIP Low-Risk $\bm{1,342(84\%)}$ $\bm{235(15\%)}$ 14 (1%)
High-Risk $\bm{42(48\%)}$ $\bm{35(40\%)}$ $\bm{11(13\%)}$
@@
Model Tight Acc. Loose Acc. ROC PRC Prec. Rec.
Orig. JHFRAT 0.69 0.98 0.88 0.51 0.63 0.29
CSO 0.79 $\bm{0.99}$ 0.93 0.66 $\bm{0.78}$ 0.44
Opt. MIP $\bm{0.82}$ 0.98 $\bm{0.94}$ $\bm{0.67}$ 0.75 $\bm{0.48}$
Figure 2: Receiver Operating Characteristic (ROC) and Precision-Recall performance curves for JHFRAT and optimized models on testing dataset. Both the CSO warm-start and optimized MIP improve over the original JHFRAT.
Figure 3 reveals how the optimized model coefficients reshape the risk score distributions compared to the original JHFRAT. The Kolmogorov-Smirnov (KS) statistic quantifies the maximum vertical separation between the cumulative distribution functions of the Safe Low-Risk and Safe High-Risk cohorts, with larger values indicating better discriminative ability. The associated p-value tests the null hypothesis that the two risk groups come from the same distribution. The original JHFRAT, CSO warm-start and optimized MIP models all have highly significant p-values (all p ¡ 0.001), confirming that all models successfully separate the risk groups, though with varying degrees of separation. The distribution of the unlabeled intervention cohort is included for reference.
Figure 3: Score distributions for all study data (test and train data together). Both CSO and MIP improve separation of high-risk and low-risk samples, at the expense of skewing high-risk and unlabeled samples to higher scores.
The overlapping distributions in the original JHFRAT model (KS = 0.620) show why current practice struggles: many true high-risk patients score in the low-to-moderate range (under-triage), while some low-risk patients score high (over-triage). This overlap drives both missed falls and unnecessary interventions. Both the CSO warm-start and optimized MIP improve on the KS statistic, at the expense of skewing the high-risk and unlabeled distributions rightward and increasing the maximum total scores from 33 for the original JHFRAT to 50 and 55 for CSO and MIP, respectively.